OpenStudy (anonymous):

(cos^2(x)-sin^2(x))/(sin(x) cos(x)) how to simplify into 2cos2x?

OpenStudy (gorv):

cos^2x-sin^2x= cos2x

OpenStudy (anonymous):

cos^2(x)-sin^2(X)=cos(2x) It's and identity 2sin(x)cos(x)=sin2(x) sin(x)cos(x)=sin(2x)/2 sooo 2cos(2x)/sin(2x)=2cot(2x) I am guessing a wrong answer?

OpenStudy (gorv):

\[\frac{ \cos2x }{ sinxcosx}\]

OpenStudy (gorv):

you r right peter well done buddy

OpenStudy (gorv):

@Auxuris got it ??

OpenStudy (anonymous):

hang on, i'm processing it

OpenStudy (gorv):

ok waitin

OpenStudy (anonymous):

but if cos^2(x)-sin^2(X)=cos(2x) then how is 2cos(2x) obtained? is it by flipping the sin(2x)/2 denominator?

OpenStudy (gorv):

if u will flip 2 will come in numerator

OpenStudy (gorv):

so 2 cos2x/sin2x

OpenStudy (anonymous):

so in other words—yes, it is by flipping the sin(2x)/2 denominator.

OpenStudy (anonymous):

thank you people :3

OpenStudy (gorv):

welcm :)